Suppose you promise to donate 2% of your company’s net income this year to local charities. If the company can manage a 4% increase over the previous year’s net income of $225,000. If your company’s net income does increase by 4%, how much will you donate?
Answer #1

This year's net income = 225,000 * (1 + 0.04) = $234,000

The amount you donate = 234,000 * 0.02 = $4,680

Can you please upvote? Thank You :-)